A small college had 2546 students in 1998, and2702 students in 2000. IF the enrollment follows a linear growth pattern, how many students will the college have in 2004? Let x be the number of years since 1998. ( In 1998, x=0)
Answer by nyc_function(2733)

Let us represent the total enrollment in any year by the general linear expression,
E = A + Bx where x is the number of years since 1998, and A, B are constants
when x=0 (ie in 1998), enrollment E = 2546
so, 2546 = A +0
so, A = 2546
in 2000, ie when x=2, enrollment was 2702
so, 2702 = 2546 +B*2
or, 2B = 2702 - 2546
or, B = 98
SO, E= 2546 +98x is the expression for enrollment
In 2004, x=6
so, Enrollment = 2546 +98*6 = 3134